Nick insists on consuming 3 times as much of y as he consumes of x (so he always has y = 3x). He will consume these goods in no other ratio. The price of x is 2 times the price of y. Nick has an endowment of 20 x’s and 75 y’s which he can trade at the going prices. He has no other source of income.

a. What is Nick’s gross demand for x? Show ALL calculations

b. What is Nick’s net demand for x? Show ALL calculations

PLEASE SHOW ALL WORK, NO EXCEL, THANK YOU!
Answer #1

Given information is,

He want to consume 3 times as much of y as x

so at any point of time,

the given price ratio,

so for solving problem easily we can just take and

Given endowments of goods are X= 20,Y=75

So his income can be written as M=20*2+75*1=115

So now the problem is much clear,he has goods with which he can earn 115,so his income is 115

prices of X and Y are 2,1 respectively

so budget line can be written as (equation 1)

his preferred consumption is always (equation2)

solve both the equations.Substitute equation 2 in equation 1.

So, Which gives

So gross demand for good X is 23.

2)Net demand for X is difference between gross demand for X and amount of X he already has.

in other words,how much of additional X (apart from what he already has) he will get by trading Y.

So net demand for X=23-20=3

PLEASE PROVIDE THE FEEDBACK IT IS AS ENCOURAGING TO ME AS TEACHER'S COMPLIMENT IS TO YOU.